OpenStudy (anonymous):

What is the expression in factored form? 9x2 – 4

OpenStudy (anonymous):

Ans: (3x-2)(3x+2)

OpenStudy (anonymous):

how did you get that.

OpenStudy (anonymous):

\[a^2-b^2= (a+b)(a-b)\] \[(3x)^2-2^2= (3x+2)(3x-2)\]